Calculate [H+] and [OH-] for each solution at 25°C
pH= 7.5
pH= 12.3
pH= 2.3


*please write in print (legibly)*
Answer #1

1) pH =7.5

pH=-log10[H+]

7.5=-log10[h+]

[H+] =antilog(-8.5)

[H+] =3.16*10^-8

PH+POh=14

POH=14-7.5=6.5

POH=-log10[OH-]

6.5=-log10[OH-]

[OH-]=antilog (-7.5)

[OH-]=3.16*10^-7

2)pH=12.3

pH=-log10[H+]

12.3=-log10[H+]

[H+]=antilog(-13.7)

[H+] =5.01*10^-13

pH+pOH=14

pOH=14-12.3=1.7

pOH=-log10[OH-]

1.7=-log10[OH-]

[OH-] =antilog (-2.3)

[OH-] =1.99*10^-2

3)ph=2.3

pH=-log10[H+]

2.3=-log10[H+]

[H+] =antilog (-3.7)

[H+] =5.01*10^-3

pH+pOH=14

pOH=14-2.3=11.7

pOH=-log10[OH-]

11.7=-log10[OH-]

[OH-]=antilog (-12.3)

[OH-]=1.99*10^-12